OpenStudy (anonymous):

Math help please! :p What is the 6th term of the geometric sequence where a1 = 625 and a3 = 25? 0.0016 0.008 0.04 0.2

OpenStudy (anonymous):

a=625 ,a3=T3=ar^2=25 Substitute a in the second equation 625*r^2=25 r^2=1/25 r=1/5 . ________ Tn=a*r^(n-1) T6=a^r(6-1) =625*1/5^(5)=1/5 .

OpenStudy (anonymous):

1/5=0.2 . Got it ?